# Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models

### v2

@@
Linear structural vector autoregressive models can be identified statistically without imposing restrictions on the model if the shocks are mutually independent and at most one of them is Gaussian. We show that this result extends to structural threshold and smooth transition vector autoregressive models incorporating a time-varying impact matrix defined as a weighted sum of the impact matrices of the regimes. We also discuss labelling of the shocks, maximum likelihood estimation of the parameters, and stationarity the model. The introduced methods are implemented to the accompanying R package sstvars. Our empirical application studies the effects of the climate policy uncertainty shock on the U.S. macroeconomy. In a structural logistic smooth transition vector autoregressive model consisting of two regimes, we find that a positive climate policy uncertainty shock decreases production in times of low economic policy uncertainty but slightly increases it in times of high economic policy uncertainty.
Keywords: smooth transition vector autoregression, nonlinear SVAR, identification by non-Gaussianity, non-Gaussian SVAR
@@ 1 Introduction
There are two main branches in the statistical identification literature: identification by heteroskedasticity (Rigobon, 2003, Lanne et al., 2010, Bacchiocchi and Fanelli, 2015, Lütkepohl and Netšunajev, 2017, Lewis, 2021, Virolainen, forthcoming, and others) and identification by non-Gaussianity (Lanne et al., 2017, Maxand, 2020, Lanne and Luoto, 2021, Lanne et al., 2023, Anttonen et al., forthcoming, and others). Under certain statistical conditions, both types of identification methods typically identify the shocks of a linear structural vector autoregressive (SVAR) model without imposing further restrictions. However, identification by heteroskedasticity without additional restrictions (as described in the cited literature) has the major drawback in nonlinear SVAR models that, for each shock, it restricts the relative magnitudes of the impact responses of the variables to stay constant over time (see the discussion in Virolainen, forthcoming, Sections 3.1 and 4). This is an undesirable property because it is generally preferable to accommodate time-variation in the (relative) impact responses. Bacchiocchi and Fanelli (2015) do propose an alternative framework for identification by heteroskedasticity that allows for variation in the impact effects of the shocks across the volatility regimes, but their method requires imposing a number of untestable zero impact effect restrictions that can be difficult to justify economically. To the best our knowledge, identifying nonlinear SVAR models by non-Gaussianity has not been studied in the previous literature.
@@
This paper contributes to the literature on identification by non-Gaussianity by extending the framework proposed by Lanne et al. (2017) to STVAR models. We also consider threshold vector autoregressive models of Tsay (1998) as special cases of STVAR models. In contrast to its linear counterpart, the impact matrix of the structural STVAR model should generally be time-varying to allow for time-variation in the impact responses of the variables to the shocks, which complicates identification. Similarly to Lanne et al. (2017), however, it turns out that identification is achieved when the shocks are mutually independent and at most one of them is Gaussian. We show that under this condition, when the impact matrix of a STVAR model is defined as a weighted sum of the impact matrices of the regimes, the shocks are readily identified up to ordering and signs. The weights of the impact matrices are the transition weights of the regimes, which we assume to be either exogenous or, for a $p$th order model, functions of the preceding $p$ observations.
In line with statistical identification literature, labelling the identified structural shocks as economic shocks requires external information. This requires labelling the same shock to the same column of the impact matrix across the regimes. In our experience, however, the identification is often weak with respect to the ordering and signs of the columns of the impact matrices in some regimes. This can be problematic because if the shocks are ordered differently in some regimes, the (generalized) impulse response functions do not estimate the effects of the intended shock but of a linear combination of structural shocks. As a solution, we propose a procedure that utilizes the almost equally well fitting local solutions with different orderings and signs of the columns of the impact matrices obtainable by our estimation method. If necessary, the identification can also be strengthened by imposing economic overidentifying short-run restrictions, which are testable.
@@
Developments related to ours include the new framework of Morioka et al. (2021) called independent innovation analysis (IIA). Morioka et al. (2021) propose using IIA for directly estimating the (conditionally) contemporaneously independent innovations of a general nonlinear vector autoregressive (VAR) model without assuming a specific functional form for the model. They show that under a certain set of (fairly general) assumptions, their method consistently estimates the independent innovations up to permutation and scalar component-wise invertible transformations. In applied macroeconometrics, the main interest is, however, typically in computing the (generalized) impulse response functions or performing other analyses that require obtaining an estimate for the parameters of the model. Estimating the parameters of a STVAR model using the framework of Morioka et al. (2021) requires first estimating the independent innovations using a specific machine learning algorithm (see Morioka et al., 2021), and then using the estimates of the innovations to obtain estimates of the parameters. However, it is not obvious how to estimate the parameters of the model from the innovations obtained by IIA, as they are identified only up to permutation and component-wise invertible transformations, and different parameter estimates would be obtained with different transformations. Moreover, Morioka et al. (2021) assume that the innovations are from a distribution in an exponential family, thereby excluding various distributions such as the popular Student’s $t$ distribution. Our approach, in turn, facilitates estimating the parameters directly by the method of maximum likelihood, and it does not exclude distributions of the independent shocks other than Gaussian distribution and those with infinite variance.
@@
Our empirical application studies the macroeconomic effects of climate policy uncertainty shocks and consider monthly U.S. data from 1987:4 to 2024:2. Following Gavriilidis et al. (2023), we measure climate policy uncertainty (CPU) with the CPU index (Gavriilidis, 2021), which is constructed based on the amount of newspaper coverage on topics related to climate policy uncertainty. We are interested in studying how the effects of the CPU shock vary depending on the level of general economic policy uncertainty. Therefore, we fit a two-regime structural logistic STVAR model using the first lag of the economic policy uncertainty index (Baker et al., 2016) as the switching variable, and allowing for smooth transitions in the intercept parameters as well as in the impact matrix. We find that a positive CPU shock decreases production in times of low economic policy uncertainty and slightly increases it in times of high economic policy uncertainty. Consumer prices, in turn, increase in response to a positive CPU shock more significantly when the level of economic policy uncertainty is high than when it is low. Our results are, hence, complementary to Gavriilidis et al. (2023), who found a positive CPU shock decreasing real activity and increasing prices, thus, transmitting to the U.S. economy as a supply shock.
@@
The rest of this paper is organized as follows. Section 2 presents the framework of reduced form STVAR models and provides examples of the covered models. Section 3 discusses identification of the shocks in structural STVAR models, presents our results on identification by non-Gaussianity, and describes a procedure for labelling the shocks in the presence of weak identification. Section 4 discusses estimation of the parameters by the method of maximum likelihood and stationarity of the model. Section 5 presents the empirical application and Section 6 concludes. Appendices provide further discussion on stationarity of the model; proofs for the stated lemmas, propositions, and theorem; and details related to the empirical application. Finally, the introduced methods have been implemented to the accompanying R package sstvars (Virolainen, 2024), which is available through the CRAN repository.
@@ 2 Reduced form STVAR models
where $\phi_{1,0},...,\phi_{M,0}\in\mathbb{R}^{d}$, $m=1,...,M$, are the intercept parameters; $A_{1,i},...,A_{M,i}\in\mathbb{R}^{d\times d}$, $i=1,...,p$, are the autoregression matrices; $u_{t}$ is a martingale difference sequence of reduced form innovations; $\Omega_{y,t}$ is the positive definite conditional covariance matrix of $u_{t}$ (conditional on $\mathcal{F}_{t-1}$), which may depend on $\alpha_{1,t},...,\alpha_{M,t}$ and $y_{t-1},...,y_{t-p}$; and $\nu$ collects any other parameters that the distribution of $u_{t}$ depends on. The transition weights $\alpha_{m,t}$ are assumed to be either nonrandom and exogenous or $\mathcal{F}_{t-1}$-measurable functions of $\{y_{t-j},j=1,...,p\}$, and they satisfy $\sum_{m=1}^{M}\alpha_{m,t}=1$ at all $t$. Thus, the types of transition weights typically used in macroeconomic applications are accommodated. The transition weights express the proportions of the regimes the process is on at each point of time and determine how the process shifts between them.
@@
It is often assumed that the coefficient matrices $A_{m,i}$, $i=1,...,p$, satisfy the usual stability condition $\det\left(I_{d}-\sum_{i=1}^{p}A_{m,i}z^{i}\right)\neq 0\ \ \text{for}\ \ |z|\leq 1,$ for each of the regimes $m=1,...,M$. However, this condition does not necessarily ensure that the model is stationary (see Kheifets and Saikkonen, 2020, and the references therein). If the transition weights are exogenous, it might be particularly challenging to establish the stationarity of the model. Assuming endogenous transition weights, we present sufficient conditions for establishing ergodicity, stationary, and mixing properties of our structural STVAR model in Appendix A (see also Section 4). Nevertheless, our main identification results discussed in Section 3 do not require stationarity of the model.
The above-described framework of STVAR models accommodates various types of transition weights employed in empirical macroeconomics. To provide a few examples, consider first the threshold VAR (TVAR) model (Tsay, 1998), which is obtained from (2.1) and (2.2) by assuming that the transition weights are defined as
@@
where $r_{0}\equiv-\infty$, $r_{M}\equiv\infty$, $m=1,...,M$, are threshold parameters, and $z_{t}$ is a switching variable. At each $t$‚ the model defined in Equations (2.1), (2.2) and (2.3) reduces to a linear VAR corresponding to one of the regimes that is determined according to the level of the switching variable $z_{t}$, which can be either exogenous or a lagged endogenous variable.
@@
where $z_{t}$ is the switching variable, $c\in\mathbb{R}$ is a location parameter, and $\gamma>0$ is a scale parameter. The location parameter $c$ determines the mid point of the transition function, i.e., the value of the switching variable when the weights are equal. The scale parameter $\gamma$, in turn, determines the smoothness of the transitions (smaller $\gamma$ implies smoother transitions), and it is assumed strictly positive so that $\alpha_{2,t}$ is increasing in $z_{t}$. The switching variable $z_{t}$ can be an exogenous variable, a lagged endogenous variable, or a linear combination of multiple such variables.
In addition specifying the transition weights or the error process $u_{t}$ in various ways, different types of models can be obtained by imposing parameter restrictions to the model defined in (2.1)-(2.2). For instance, a model that is linear in the autoregressive (AR) coefficients is obtained by assuming that they are constants across the regimes, i.e., $A_{m,i}=A_{i}$ and $\phi_{m,0}=\phi_{0}$ for all $i=1,...,p$ and $M=1,...,M$. The model then simplifies to
@@
Imposing constant AR matrices across the regimes but allowing the intercepts to vary in time is a particularly useful special case of (2.1)-(2.2), as constant AR matrices substantially decrease the number of parameters compared to the fully nonlinear specification. However, the model still allows for a time-varying conditional mean via time-varying intercepts as well as shifts in the volatility regime through the error process $u_{t}$. This specification is obtained from (2.1)-(2.2) by assuming $A_{m,i}=A_{i}$ for all $i=1,...,p$ and $M=1,...,M$, simplifying the model to:
$y_{t}=\sum_{m=1}^{M}\alpha_{m,t}\phi_{m,0}+\sum_{i=1}^{p}A_{i}y_{t-i}+u_{t},\quad u_{t}\sim MD(0,\Omega_{y,t},\nu).$ (2.6)
Finally, note that while we focus on TVAR and STVAR models here for concreteness and because these models are implemented to the accompanying R package sstvars (Virolainen, 2024), the identification results of this paper directly extend to many other nonlinear SVAR models as well. This includes models incorporating discrete regime switches such as Markov-Switching VAR models (Krolzig, 1997) and mixture VAR models (e.g., Fong et al., 2007), and more generally models whose impact matrix is defined as a weighted sum of the impact matrices of the regimes and the time $t$ weights are known at the time $t$.
@@ 3.1 Structural STVAR models
where $B_{y,t}$ is an invertible ($d\times d$) impact matrix that governs the contemporaneous relationships of the shocks and may depend on $\alpha_{1,t},...,\alpha_{M,t}$ and $y_{t-1},...,y_{t-p}$. In other words, assuming a unit variance normalization for the structural shocks, the identification problem amounts to finding an impact matrix $B_{y,t}$ such that $\Omega_{y,t}=B_{y,t}B_{y,t}^{\prime}$. The impact matrix is not, however, generally uniquely identified without imposing further restrictions on the model, since there exists many such decompositions of $\Omega_{y,t}$. Various types of restrictions that recover the structural shocks have been proposed in the literature. Many of them are discussed in Ramey (2016) and Kilian and Lütkepohl (2017), for example.
Following the literature on statistical identification by non-Gaussianity in the framework proposed by Lanne et al. (2017), our identifying restriction is that the shocks are mutually independent and that at most one of them is Gaussian. With Gaussian shocks, different decompositions of the conditional covariance matrix $\Omega_{y,t}=B_{y,t}B_{y,t}^{\prime}$ lead to observationally equivalent models. However, when the shocks are mutually independent and at most of them is Gaussian, different decompositions generally lead to models that are not observationally equivalent, as shown in Lanne et al. (2017) for a static impact matrix. We extend this result to a time-varying impact matrix, and therefore, it is useful to parametrize the structural model directly with the impact matrix $B_{y,t}$.
@@
Since we incorporate time-variation in the impact matrix, its functional form needs to be specified. A natural specification for a STVAR model is to assume a constant impact matrix for each of the regimes and define the impact matrix of the process to be their weighted sum as
@@
The result of Lemma 1 holds almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, which means that the set where the matrices $B_{1},...,B_{M}$ that are such that $B_{y,t}$ is singular for some $t$ has a Lebesgue measure zero. This implies that invertible matrices $B_{1},...,B_{M}$ generally lead to an invertible impact matrix $B_{y,t}$, excluding some special cases of $B_{1},...,B_{M}$ for which this result does not hold (e.g., if $B_{2}=-B_{1}$, then $B_{y,t}=0$ when $\alpha_{1,t}=\alpha_{2,t}=0.5$). To further improve intuition, observe that the result of Lemma 1 implies that if the matrices $B_{1},...,B_{M}$ are drawn by random from a continuous distribution, the probability of obtaining matrices such that $B_{y,t}$ is singular for some $t$ is zero. In practice, the parameter space should be simply restricted in the estimation so that parameter values leading to an impact matrix $B_{t}$ that is singular for some $t=1,...,T$ are excluded.
@@
where $\Omega_{m}\equiv B_{m}B_{m}^{\prime}$ and $\Omega_{m,n}\equiv B_{m}B_{n}^{\prime}$. The conditional covariance matrix of $y_{t}$ can be thereby described as a weighted sum of $M^{2}$ matrices with the weights varying in time according to the transition weights $\alpha_{m,t}$, $m=1,...,M$. To facilitate the interpretation of the conditional covariance matrix (3.4), suppose the process is completely in one of the regimes at some $t$, i.e., $\alpha_{\tilde{m},t}=1$ for some $\tilde{m}\in\{1,...,M\}$ and $\alpha_{m,t}=0$ for $m\neq\tilde{m}$. Then, $\Omega_{y,t}$ reduces to $B_{\tilde{m}}B_{\tilde{m}}^{\prime}=\Omega_{\tilde{m}}$, implying that this constitutes as the conditional covariance matrix of the regime $\tilde{m}$. When the process is not completely in any of the regimes, $\alpha_{\tilde{m},t}\in(0,1)$ for some regime $\tilde{m}$$\Omega_{y,t}$ depends both on the conditional covariance matrices of the regimes $\Omega_{1},...,\Omega_{M}$ and on the cross terms $\Omega_{m,n}$. Hence, this specification of the structural model is different to the conventional reduced form STVAR specification in which the conditional covariance matrix of $y_{t}$ has the form $B_{y,t}B_{y,t}^{\prime}=\sum_{m=1}^{M}\alpha_{m,t}\Omega_{m}$ for the covariance matrices $\Omega_{1},...,\Omega_{M}$ of the regimes.
Specifying the conditional covariance matrix as $\Omega_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}\Omega_{m}$ would either require parametrizing the model directly with $\Omega_{y,t}$ or imposing some specific structure on $B_{y,t}$ that would ensure that this identity holds at all $t$. In the former case, identification of the shocks requires further restrictions to be imposed on the model (even when the shocks are non-Gaussian) as discussed at the beginning of this section. In the latter case, the restrictions are implicitly imposed by the specific form of the impact matrix ensuring that $B_{y,t}B_{y,t}=\sum_{m=1}^{M}\alpha_{m,t}\Omega_{m}$ for all $t$. One of such parametrizations of $B_{y,t}$ aligns with the identification by heteroskedasticity (Lütkepohl and Netšunajev, 2017), which restricts the relative impact responses of the variables to each shock time-invariant (see also the related discussion in Virolainen, forthcoming). In either case, the identifying restrictions may be challenging to justify economically. Our specification of $\Omega_{y,t}$ in (3.3) and (3.4), in turn, facilitates exploiting non-Gaussianity of the shocks in their identification, as we show the next section.

### v3

@@
Linear structural vector autoregressive models can be identified statistically without imposing restrictions on the model if the shocks are mutually independent and at most one of them is Gaussian. We show that this result extends to structural threshold and smooth transition vector autoregressive models incorporating a time-varying impact matrix defined as a weighted sum of the impact matrices of the regimes. We also discuss the problem of labelling the shocks, estimation of the parameters, and stationarity the model. The introduced methods are implemented to the accompanying R package sstvars. Our empirical application studies the effects of the climate policy uncertainty shock on the U.S. macroeconomy. In a structural logistic smooth transition vector autoregressive model consisting of two regimes, we find that a positive climate policy uncertainty shock decreases production and increases inflation in times of both low and high economic policy uncertainty, but its inflationary effects are stronger in the periods of high economic policy uncertainty.
Keywords: smooth transition vector autoregression, threshold VAR, nonlinear SVAR, identification by non-Gaussianity, non-Gaussian SVAR
@@ 1 Introduction
There are two main branches in the statistical identification literature: identification by heteroscedasticity (Rigobon, 2003, Lanne et al., 2010, Bacchiocchi and Fanelli, 2015, Lütkepohl and Netšunajev, 2017, Lewis, 2021, Virolainen, 2025b, and others) and identification by non-Gaussianity (Lanne et al., 2017, Lanne and Luoto, 2021, Lanne et al., 2023, and others). Under certain statistical conditions, both types of identification methods typically identify the shocks of a linear structural vector autoregressive (SVAR) model without imposing further restrictions. However, identification by heteroscedasticity without additional restrictions (as described in the cited literature) has the major drawback in nonlinear SVAR models that, for each shock, it restricts the relative magnitudes of the impact responses of the variables to stay constant over time (see, e.g., Lütkepohl and Netšunajev, 2017). This is an undesirable property because it is generally preferable to accommodate time-variation in the (relative) impact responses. To the best our knowledge, identifying nonlinear SVAR models by non-Gaussianity has not been studied in the previous literature.
@@
This paper contributes to the literature on identification by non-Gaussianity by extending the framework proposed by Lanne et al. (2017) to STVAR models. We also consider threshold vector autoregressive models of Tsay (1998) as special cases of STVAR models. In contrast to its linear counterpart, the impact matrix of the structural STVAR model should generally be time-varying to allow for time-variation in the impact responses of the variables to the shocks, which complicates identification. Nonetheless, similarly to Lanne et al. (2017), it turns out that identification is achieved when the shocks are mutually independent and at most one of them is Gaussian. We show that under this condition, when the impact matrix of a STVAR model is defined as a weighted sum of the impact matrices of the regimes, the shocks are readily identified up to ordering and signs. The weights of the impact matrices are the transition weights of the regimes, which we assume to be either exogenous, logistic, or of the threshold form, but our results can be generalized to other suitable weight functions as well. In line with statistical identification literature, however, external information is required to labelling the identified structural shocks as economic shocks. Since the identification can, in our experience, be also weak with respect to the ordering and signs of the columns of the regime-specific impact matrices, we recommend adopting a blended identification strategy that combines identification by non-Gaussianity with additional information (cf. Carriero et al., 2024).
@@
Developments related to ours include the new framework of Morioka et al. (2021) called independent innovation analysis (IIA). The IIA estimates the (conditionally) contemporaneously independent innovations of a nonlinear SVAR model without imposing any specific functional form. Under fairly general assumptions, their method consistently estimates the independent innovations up to permutation and scalar component-wise invertible transformations. While the IIA does seem promising, it is not obvious how to apply it in empirical macroeconometrics, as identification only up to such transformations is, in our understanding, insufficient for structural analysis. Moreover, Morioka et al. (2021) assume innovations follow a distribution in the exponential family, thereby excluding various alternatives. In contrast, our approach facilitates estimation of the parameters of an identified nonlinear SVAR model, enabling structural analysis without restricting the distribution of independent shocks beyond ruling out Gaussian and infinite-variance cases.
@@
Our empirical application studies the macroeconomic effects of climate policy uncertainty shocks and consider monthly U.S. data from 1987:4 to 2024:12. Following Khalil and Strobel (2023) and Huang and Punzi (2024), we measure climate policy uncertainty (CPU) with the CPU index (Gavriilidis, 2021), which is constructed based on the amount of newspaper coverage on topics related to CPU. We are interested in studying how the effects of the CPU shock vary depending on the level of economic policy uncertainty (EPU). Therefore, we fit a two-regime structural logistic STVAR model using the first lag of the EPU index (Baker et al., 2016) as the switching variable. We find that a positive CPU shock decreases production and increases inflation in times both low and high EPU, but its inflationary effects are stronger in the periods of high EPU. Our results are, hence, in line with the previous literature suggesting that a positive CPU shock decreases production (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024) and increases inflation (Khalil and Strobel, 2023, and Huang and Punzi, 2024).
@@
The rest of this paper is organized as follows. Section 2 presents the framework of reduced form STVAR models and provides examples of the covered models. Section 3 discusses identification of the shocks in structural STVAR models, presents our results on identification by non-Gaussianity, and discusses the problem of labelling the shocks. Section 4 discusses stationarity of the model and proposes estimating its parameters with a penalized likelihood -based estimator using a three-phase procedure. Section 5 presents the empirical application and Section 6 concludes. Appendices provide detailed discussion on ergodic stationarity of the model; proofs for the stated lemmas, propositions, and theorem; details related to the proposed estimation procedure; a Monte Carlo study assessing the performance of the proposed estimator; and details related to the empirical application. Finally, the introduced methods have been implemented to the accompanying R package sstvars (Virolainen, 2025a), which is available through the CRAN repository.
@@ 2 Reduced form STVAR models
where $\phi_{1,0},...,\phi_{M,0}\in\mathbb{R}^{d}$, $m=1,...,M$, are the intercept parameters; $A_{1,i},...,A_{M,i}\in\mathbb{R}^{d\times d}$, $i=1,...,p$, are the autoregression matrices; $u_{t}$ is a martingale difference sequence of reduced form innovations; $\Omega_{y,t}$ is the positive definite conditional covariance matrix of $u_{t}$ (conditional on $\mathcal{F}_{t-1}$), which may depend on $\alpha_{1,t},...,\alpha_{M,t}$ and $y_{t-1},...,y_{t-p}$; and $\nu$ collects any other parameters that the distribution of $u_{t}$ depends on. The transition weights $\alpha_{m,t}$ are assumed to be either exogenous (nonrandom) or $\mathcal{F}_{t-1}$-measurable functions of $\{y_{t-j},j=1,...,p\}$, and to satisfy $\sum_{m=1}^{M}\alpha_{m,t}=1$ at all $t$. Thus, the types of transition weights typically used in macroeconomic applications are accommodated. The transition weights express the proportions of the regimes the process is on at each point of time and determine how the process shifts between them.
@@
Our framework of STVAR models accommodates various types of transition weights employed in empirical macroeconomics. To provide a few examples, consider first the threshold VAR (TVAR) model (Tsay, 1998), which is obtained from (2.1) and (2.2) by assuming that the transition weights are defined as
@@
where $r_{0}\equiv-\infty$, $r_{M}\equiv\infty$, $m=1,...,M$, are threshold parameters, and $z_{t}$ is a switching variable. At each $t$‚ the model defined in Equations (2.1), (2.2) and (2.3) reduces to a linear VAR corresponding to one of the regimes that is determined according to the level of the switching variable $z_{t}$, which can be an exogenous variable, a lagged endogenous variable, or a linear combination of multiple such variables, for example.
@@
where $z_{t}$ is the switching variable, $c\in\mathbb{R}$ is a location parameter, and $\gamma>0$ is a scale parameter. The location parameter $c$ determines the midpoint of the transition function, i.e., the value of the switching variable when the weights are equal. The scale parameter $\gamma$, in turn, determines the smoothness of the transitions (smaller $\gamma$ implies smoother transitions), and it is assumed strictly positive so that $\alpha_{2,t}$ is increasing in $z_{t}$.
In addition to specifying the transition weights or the error process $u_{t}$ in various ways, different types of models can be obtained by imposing parameter restrictions to the model defined in (2.1)-(2.2). For instance, a model that is linear in the autoregressive (AR) coefficients is obtained by assuming that they are constants across the regimes, i.e., $A_{m,i}=A_{i}$ and $\phi_{m,0}=\phi_{0}$ for all $i=1,...,p$ and $M=1,...,M$. The model then simplifies to
@@ 3.1 Structural STVAR models
where $B_{y,t}$ is an invertible ($d\times d$) impact matrix that governs the contemporaneous relationships of the shocks and may depend on $\alpha_{1,t},...,\alpha_{M,t}$ and $y_{t-1},...,y_{t-p}$. In other words, assuming a unit variance normalization for the structural shocks, the identification problem amounts to finding an impact matrix $B_{y,t}$ such that $\Omega_{y,t}=B_{y,t}B_{y,t}^{\prime}$. The impact matrix is not, however, generally uniquely identified without imposing further restrictions on the model, since there exists many such decompositions of $\Omega_{y,t}$. Various types of restrictions that recover the structural shocks have been proposed in the literature, many of which are discussed in Kilian and Lütkepohl (2017), for example.
Following the literature on statistical identification by non-Gaussianity in the framework proposed by Lanne et al. (2017), our identifying restriction is that the shocks are mutually independent and that at most one of them is Gaussian. Under this assumption, unlike in the Gaussian case, different decompositions of the conditional covariance matrix $\Omega_{y,t}=B_{y,t}B_{y,t}^{\prime}$ generally lead to observationally distinct models, effectively identifying a static impact matrix and the shocks (Lanne et al., 2017). We extend this result to a time-varying impact matrix, and therefore, it is useful to parametrize the structural model directly with $B_{y,t}$.
@@
To incorporate time-variation in the impact matrix, we specify its functional form. A natural specification for STVAR models is to assume a constant impact matrix for each of the regimes and define the impact matrix of the process to be their weighted sum as
@@
The result of Lemma 1 holds almost everywhere in $[B_{1}:...:B_{M}]\in\mathbb{R}^{d\times dM}$, which means that the set where the matrices $B_{1},...,B_{M}$ are such that $B_{y,t}$ is singular for some $t$ has a Lebesgue measure zero. This implies that invertible matrices $B_{1},...,B_{M}$ generally lead to an invertible impact matrix $B_{y,t}$, excluding some special cases of $B_{1},...,B_{M}$ for which this result does not hold (e.g., if $B_{2}=-B_{1}$, then $B_{y,t}=0$ when $\alpha_{1,t}=\alpha_{2,t}=0.5$). To further improve intuition, note that the result of Lemma 1 implies that if the matrices $B_{1},...,B_{M}$ are drawn by random from a continuous distribution, the probability of obtaining matrices such that $B_{y,t}$ is singular for some $t$ is zero. In practice, the parameter space should be simply restricted in the estimation so that parameter values leading to $B_{y,t}$ that is singular for some $t$ are excluded.
@@
where $\Omega_{m}\equiv B_{m}B_{m}^{\prime}$ and $\Omega_{m,n}\equiv B_{m}B_{n}^{\prime}$. The conditional covariance matrix of $y_{t}$ can be thus described as a weighted sum of $M^{2}$ matrices with the weights varying in time according to the transition weights $\alpha_{m,t}$, $m=1,...,M$. If the process is completely in one of the regimes, i.e., $\alpha_{\tilde{m},t}=1$ for some $\tilde{m}\in\{1,...,M\}$, then $\Omega_{y,t}$ reduces to $\Omega_{\tilde{m}}$, implying that this constitutes as the conditional covariance matrix of Regime $\tilde{m}$. When the process is not completely in any of the regimes, $\Omega_{y,t}$ depends both on the regime-specific covariance matrices and the cross terms. Hence, this specification of the structural model is different to the conventional reduced form STVAR specification in which the conditional covariance matrix of $y_{t}$ is a weighted sum of the covariance matrices of the regimes. Nonetheless, our model specification facilitates exploiting non-Gaussianity of the shocks in their identification, as we show in the next section.